\begin{thebibliography}{41}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Halbert and French(1957)]{PhysRev.105.1563}
E.~C. Halbert and J.~B. French.
\newblock Shell model for the positive-parity states of $^{15}${N}.
\newblock \emph{Phys. Rev.}, 105:\penalty0 1563--1569, Mar 1957.

\bibitem[Forss\'en et~al.(2018)Forss\'en, Carlsson, Johansson, S\"a\"af,
  Bansal, Hagen, and Papenbrock]{PhysRevC.97.034328}
C.~Forss\'en, B.~D. Carlsson, H.~T. Johansson, D.~S\"a\"af, A.~Bansal,
  G.~Hagen, and T.~Papenbrock.
\newblock Large-scale exact diagonalizations reveal low-momentum scales of
  nuclei.
\newblock \emph{Phys. Rev. C}, 97:\penalty0 034328, 2018.

\bibitem[Suzuki et~al.(2006)Suzuki, Chiba, Yoshida, Kajino, and
  Otsuka]{PhysRevC.74.034307}
Toshio Suzuki, Satoshi Chiba, Takashi Yoshida, Toshitaka Kajino, and Takaharu
  Otsuka.
\newblock Neutrino-nucleus reactions based on new shell model hamiltonians.
\newblock \emph{Phys. Rev. C}, 74:\penalty0 034307, 2006.

\bibitem[Anand et~al.(2014)Anand, Fitzpatrick, and Haxton]{PhysRevC.89.065501}
Nikhil Anand, A.~Liam Fitzpatrick, and W.~C. Haxton.
\newblock Weakly interacting massive particle-nucleus elastic scattering
  response.
\newblock \emph{Phys. Rev. C}, 89:\penalty0 065501, 2014.

\bibitem[Horoi and Brown(2013)]{PhysRevLett.110.222502}
M.~Horoi and B.~A. Brown.
\newblock Shell-model analysis of the $^{136}\mathrm{Xe}$ double beta decay
  nuclear matrix elements.
\newblock \emph{Phys. Rev. Lett.}, 110:\penalty0 222502, May 2013.

\bibitem[Haxton and Wieman(2001)]{haxton2001atomic}
WC~Haxton and Carl~E Wieman.
\newblock Atomic parity nonconservation and nuclear anapole moments.
\newblock \emph{Annual Review of Nuclear and Particle Science}, 51\penalty0
  (1):\penalty0 261--293, 2001.

\bibitem[Brussard and Glaudemans(1977)]{BG77}
P.J. Brussard and P.W.M. Glaudemans.
\newblock \emph{Shell-model applications in nuclear spectroscopy}.
\newblock North-Holland Publishing Company, Amsterdam, 1977.

\bibitem[Brown and Wildenthal(1988)]{br88}
B.~A. Brown and B.~H. Wildenthal.
\newblock Status of the nuclear shell model.
\newblock \emph{{Annual Review of Nuclear and Particle Science}}, 38:\penalty0
  29--66, 1988.

\bibitem[Caurier et~al.(2005)Caurier, Martinez-Pinedo, Nowacki, Poves, and
  Zuker]{ca05}
E.~Caurier, G.~Martinez-Pinedo, F.~Nowacki, A.~Poves, and A.~P. Zuker.
\newblock The shell model as a unified view of nuclear structure.
\newblock \emph{{Reviews of Modern Physics}}, 77:\penalty0 427--488, 2005.

\bibitem[Brown and Richter(2006)]{PhysRevC.74.034315}
B.~Alex Brown and W.~A. Richter.
\newblock New ``{USD}'' hamiltonians for the $\mathit{sd}$ shell.
\newblock \emph{Phys. Rev. C}, 74:\penalty0 034315, 2006.

\bibitem[Entem and Machleidt(2003)]{PhysRevC.68.041001}
D.~R. Entem and R.~Machleidt.
\newblock Accurate charge-dependent nucleon-nucleon potential at fourth order
  of chiral perturbation theory.
\newblock \emph{Phys. Rev. C}, 68:\penalty0 041001, 2003.

\bibitem[Wiringa et~al.(1995)Wiringa, Stoks, and
  Schiavilla]{wiringa1995accurate}
Robert~B Wiringa, VGJ Stoks, and R~Schiavilla.
\newblock Accurate nucleon-nucleon potential with charge-independence breaking.
\newblock \emph{Physical Review C}, 51\penalty0 (1):\penalty0 38, 1995.

\bibitem[Shirokov et~al.(2016)Shirokov, Shin, Kim, Sosonkina, Maris, and
  Vary]{shirokov2016n3lo}
AM~Shirokov, IJ~Shin, Y~Kim, M~Sosonkina, P~Maris, and JP~Vary.
\newblock {N3LO} {NN} interaction adjusted to light nuclei in ab exitu
  approach.
\newblock \emph{Physics Letters B}, 761:\penalty0 87--91, 2016.

\bibitem[Dyhdalo et~al.(2016)Dyhdalo, Furnstahl, Hebeler, and
  Tews]{PhysRevC.94.034001}
A.~Dyhdalo, R.~J. Furnstahl, K.~Hebeler, and I.~Tews.
\newblock Regulator artifacts in uniform matter for chiral interactions.
\newblock \emph{Phys. Rev. C}, 94:\penalty0 034001, 2016.

\bibitem[Navr{\'a}til et~al.(2009)Navr{\'a}til, Quaglioni, Stetcu, and
  Barrett]{navratil2009recent}
Petr Navr{\'a}til, Sofia Quaglioni, Ionel Stetcu, and Bruce~R Barrett.
\newblock Recent developments in no-core shell-model calculations.
\newblock \emph{Journal of Physics G: Nuclear and Particle Physics},
  36\penalty0 (8):\penalty0 083101, 2009.

\bibitem[Barrett et~al.(2013)Barrett, Navr{\'a}til, and Vary]{barrett2013ab}
Bruce~R Barrett, Petr Navr{\'a}til, and James~P Vary.
\newblock Ab initio no core shell model.
\newblock \emph{Progress in Particle and Nuclear Physics}, 69:\penalty0
  131--181, 2013.

\bibitem[Dikmen et~al.(2015)Dikmen, Lisetskiy, Barrett, Maris, Shirokov, and
  Vary]{PhysRevC.91.064301}
E.~Dikmen, A.~F. Lisetskiy, B.~R. Barrett, P.~Maris, A.~M. Shirokov, and J.~P.
  Vary.
\newblock Ab initio effective interactions for $sd$-shell valence nucleons.
\newblock \emph{Phys. Rev. C}, 91:\penalty0 064301, 2015.

\bibitem[Jansen et~al.(2014)Jansen, Engel, Hagen, Navratil, and
  Signoracci]{PhysRevLett.113.142502}
G.~R. Jansen, J.~Engel, G.~Hagen, P.~Navratil, and A.~Signoracci.
\newblock Ab initio coupled-cluster effective interactions for the shell model:
  Application to neutron-rich oxygen and carbon isotopes.
\newblock \emph{Phys. Rev. Lett.}, 113:\penalty0 142502, 2014.

\bibitem[Stroberg et~al.(2017)Stroberg, Calci, Hergert, Holt, Bogner, Roth, and
  Schwenk]{PhysRevLett.118.032502}
S.~R. Stroberg, A.~Calci, H.~Hergert, J.~D. Holt, S.~K. Bogner, R.~Roth, and
  A.~Schwenk.
\newblock Nucleus-dependent valence-space approach to nuclear structure.
\newblock \emph{Phys. Rev. Lett.}, 118:\penalty0 032502, 2017.

\bibitem[Whitehead et~al.(1977)Whitehead, Watt, Cole, and Morrison]{Lanczos}
R.~R. Whitehead, A.~Watt, B.~J. Cole, and I.~Morrison.
\newblock Computational methods for shell model calculations.
\newblock \emph{{Advances in Nuclear Physics}}, 9:\penalty0 123--176, 1977.

\bibitem[Shao et~al.(2018)Shao, Aktulga, Yang, Ng, Maris, and
  Vary]{shao2018accelerating}
Meiyue Shao, H~Metin Aktulga, Chao Yang, Esmond~G Ng, Pieter Maris, and James~P
  Vary.
\newblock Accelerating nuclear configuration interaction calculations through a
  preconditioned block iterative eigensolver.
\newblock \emph{Computer Physics Communications}, 222:\penalty0 1--13, 2018.

\bibitem[Caurier and Nowacki(1999)]{ANTOINE}
E.~Caurier and F~Nowacki.
\newblock Present status of shell model techniques.
\newblock \emph{{Scopus Preview}}, 30:\penalty0 705--714, 1999.

\bibitem[Sternberg et~al.(2008)Sternberg, Ng, Yang, Maris, Vary, Sosonkina, and
  Le]{MFDn}
P.~Sternberg, E.~Ng, C.~Yang, P.~Maris, J.P. Vary, M.~Sosonkina, and H.~Viet
  Le.
\newblock {Accelerating configuration interaction calculations for nuclear
  structure}.
\newblock \emph{The Proceedings of the 2008 ACM/IEEE Conference on
  Supercomputing}, 2008.

\bibitem[Johnson et~al.(2013)Johnson, Ormand, and Krastev]{BIGSTICK}
Calvin~W. Johnson, W.~Erich Ormand, and Plamen~G. Krastev.
\newblock Factorization in large-scale many-body calculations.
\newblock \emph{Computer Physics Communications}, 184:\penalty0 2761--2774,
  2013.

\bibitem[Johnson et~al.(2018)Johnson, Ormand, McElvain, and
  Shan]{johnson2018bigstick}
Calvin~W Johnson, W~Erich Ormand, Kenneth~S McElvain, and Hongzhang Shan.
\newblock Bigstick: A flexible configuration-interaction shell-model code.
\newblock \emph{arXiv preprint arXiv:1801.08432}, 2018.

\bibitem[Shimizu(2013)]{shimizu2013nuclear}
Noritaka Shimizu.
\newblock Nuclear shell-model code for massive parallel computation, {KSHELL}.
\newblock \emph{arXiv preprint arXiv:1310.5431}, 2013.

\bibitem[Brown et~al.(1985)Brown, Etchegoyen, and Rae]{OXBASH}
B.A. Brown, A.~Etchegoyen, and W.D.M. Rae.
\newblock {Computer code OXBASH: the Oxford University-Buenos Aires-MSU shell
  model code}.
\newblock \emph{Michigan State University Cyclotron Laboratory Report No. 524},
  1985.

\bibitem[Brown and Rae(2014)]{NuShellX}
B.~A. Brown and W.~D.~M. Rae.
\newblock {The Shell-Model Code NuShellX@MSU}.
\newblock \emph{{Nuclear Data Sheets}}, 120:\penalty0 115--118, 2014.

\bibitem[Dytrych et~al.(2013)Dytrych, Launey, Draayer, Maris, Vary, Saule,
  Catalyurek, Sosonkina, Langr, and Caprio]{PhysRevLett.111.252501}
T.~Dytrych, K.~D. Launey, J.~P. Draayer, P.~Maris, J.~P. Vary, E.~Saule,
  U.~Catalyurek, M.~Sosonkina, D.~Langr, and M.~A. Caprio.
\newblock Collective modes in light nuclei from first principles.
\newblock \emph{Phys. Rev. Lett.}, 111:\penalty0 252501, 2013.

\bibitem[McCoy et~al.(2018)McCoy, Caprio, and Dytrych]{mccoysymplectic}
Anna~E McCoy, Mark~A Caprio, and Tomas Dytrych.
\newblock Symplectic no-core configuration interaction framework for ab initio
  nuclear structure.
\newblock \emph{Ann. Acad. Rom. Sci. Ser. Chem. Phys. Sci.}, 3:\penalty0 17,
  2018.

\bibitem[Dytrych et~al.(2016)Dytrych, Maris, Launey, Draayer, Vary, Langr,
  Saule, Caprio, Catalyurek, and Sosonkina]{dytrych2016efficacy}
Tom{\'a}s Dytrych, Pieter Maris, Kristina~D Launey, Jerry~P Draayer, James~P
  Vary, Daniel Langr, Erik Saule, MA~Caprio, U~Catalyurek, and Masha Sosonkina.
\newblock Efficacy of the {SU}(3) scheme for ab initio large-scale calculations
  beyond the lightest nuclei.
\newblock \emph{Computer Physics Communications}, 207:\penalty0 202--210, 2016.

\bibitem[Caprio et~al.(2012)Caprio, Maris, and Vary]{PhysRevC.86.034312}
M.~A. Caprio, P.~Maris, and J.~P. Vary.
\newblock Coulomb-sturmian basis for the nuclear many-body problem.
\newblock \emph{Phys. Rev. C}, 86:\penalty0 034312, 2012.

\bibitem[Constantinou et~al.(2017)Constantinou, Caprio, Vary, and
  Maris]{constantinou2017natural}
Chrysovalantis Constantinou, Mark~A Caprio, James~P Vary, and Pieter Maris.
\newblock Natural orbital description of the halo nucleus 6 he.
\newblock \emph{Nuclear Science and Techniques}, 28\penalty0 (12):\penalty0
  179, 2017.

\bibitem[Heng et~al.(2017)Heng, Vary, and Maris]{heng2017ab}
Taihua Heng, James~P Vary, and Pieter Maris.
\newblock Ab initio no-core properties of li 7 and be 7 with the jisp16 and
  chiral nnlo opt interactions.
\newblock \emph{Physical Review C}, 95\penalty0 (1):\penalty0 014306, 2017.

\bibitem[Horoi et~al.(1999)Horoi, Volya, and Zelevinsky]{PhysRevLett.82.2064}
Mihai Horoi, Alexander Volya, and Vladimir Zelevinsky.
\newblock Chaotic wave functions and exponential convergence of low-lying
  energy eigenvalues.
\newblock \emph{Phys. Rev. Lett.}, 82:\penalty0 2064--2067, 1999.

\bibitem[Coon et~al.(2012)Coon, Avetian, Kruse, van Kolck, Maris, and
  Vary]{PhysRevC.86.054002}
S.~A. Coon, M.~I. Avetian, M.~K.~G. Kruse, U.~van Kolck, P.~Maris, and J.~P.
  Vary.
\newblock Convergence properties of ab initio calculations of light nuclei in a
  harmonic oscillator basis.
\newblock \emph{Phys. Rev. C}, 86:\penalty0 054002, 2012.

\bibitem[More et~al.(2013)More, Ekstr\"om, Furnstahl, Hagen, and
  Papenbrock]{PhysRevC.87.044326}
S.~N. More, A.~Ekstr\"om, R.~J. Furnstahl, G.~Hagen, and T.~Papenbrock.
\newblock Universal properties of infrared oscillator basis extrapolations.
\newblock \emph{Phys. Rev. C}, 87:\penalty0 044326, 2013.

\bibitem[Wendt et~al.(2015)Wendt, Forss\'en, Papenbrock, and
  S\"a\"af]{PhysRevC.91.061301}
K.~A. Wendt, C.~Forss\'en, T.~Papenbrock, and D.~S\"a\"af.
\newblock Infrared length scale and extrapolations for the no-core shell model.
\newblock \emph{Phys. Rev. C}, 91:\penalty0 061301, 2015.

\bibitem[Furnstahl et~al.(2012)Furnstahl, Hagen, and
  Papenbrock]{PhysRevC.86.031301}
R.~J. Furnstahl, G.~Hagen, and T.~Papenbrock.
\newblock Corrections to nuclear energies and radii in finite oscillator
  spaces.
\newblock \emph{Phys. Rev. C}, 86:\penalty0 031301, 2012.

\bibitem[Hagen et~al.(2010)Hagen, Papenbrock, Dean, and
  Hjorth-Jensen]{hagen2010ab}
G~Hagen, T~Papenbrock, David~J Dean, and Morten Hjorth-Jensen.
\newblock Ab initio coupled-cluster approach to nuclear structure with modern
  nucleon-nucleon interactions.
\newblock \emph{Physical Review C}, 82\penalty0 (3):\penalty0 034330, 2010.

\bibitem[Negoita et~al.(2018)Negoita, Luecke, Vary, Maris, Shirokov, Shin, Kim,
  Ng, and Yang]{negoita2018deep}
Gianina~Alina Negoita, Glenn~R Luecke, James~P Vary, Pieter Maris, Andrey~M
  Shirokov, Ik~Jae Shin, Youngman Kim, Esmond~G Ng, and Chao Yang.
\newblock Deep learning: A tool for computational nuclear physics.
\newblock \emph{arXiv preprint arXiv:1803.03215}, 2018.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{graphicx}
\usepackage{natbib}

\newcommand\pubnumber{}
\newcommand\pubdate{\today}

\def\Title#1{\begin{center} {\Large #1 } \end{center}}
\def\Author#1{\begin{center}{ \sc #1} \end{center}}
\def\Address#1{\begin{center}{ \it #1} \end{center}}

\newcommand\pubblock{\rightline{\begin{tabular}{l} \pubnumber\\
         \pubdate  \end{tabular}}}
\newenvironment{Abstract}{\begin{quotation}  }{\end{quotation}}
\newenvironment{Presented}{\begin{quotation} \begin{center} 
             PRESENTED AT\end{center}\bigskip 
      \begin{center}\begin{large}}{\end{large}\end{center} \end{quotation}}
\def\Acknowledgements{\bigskip  \bigskip \begin{center} \begin{large}
             \bf ACKNOWLEDGEMENTS \end{large}\end{center}}

\def\beq{\begin{equation}}
\def\eeq#1{\label{#1}\end{equation}}
\def\eeqn{\end{equation}}

\def\beqa{\begin{eqnarray}}
\def\eeqa#1{\label{#1}\end{eqnarray}}
\def\eeqan{\end{eqnarray}}

\let\bar=\overbar

\def\Dslash{\not{\hbox{\kern-4pt $D$}}}
\def\dslash{\not{\hbox{\kern-2pt $\del$}}}

\def\msb{{\bar{\ssstyle M \kern -1pt S}}}

\begin{document}
\begin{titlepage}
\pubblock

\vfill
\Title{Current Status of Very-Large-Basis Hamiltonian Diagonalizations for Nuclear Physics}
\vfill
\Author{Calvin W. Johnson\footnote{This material is based upon work supported by the U.S. Department of Energy, 
Office of Science, Office of Nuclear Physics, under Award Number  DE-FG02-03ER41272}}
\Address{Department of Physics\\
San Diego State University, 92182-1233, CA, USA}
\vfill
\begin{Abstract}
Today there are a plethora of many-body techniques for calculating nuclear wave functions and matrix elements. I review the status of that reliable workhorse, the interacting shell model, a.k.a. configuration-interaction methods, a.k.a. Hamiltonian diagonalization, and survey its advantages and disadvantages.  With modern supercomputers one can tackle dimensions up to about 20 billion! I  discuss how we got 
there and where we might go in the near future.
\end{Abstract}
\vfill
\begin{Presented}
Thirteenth Conference on the Intersections of Particle and Nuclear Physics\\
Indian Wells, CA,  May 28--June 3, 2018
\end{Presented}
\vfill
\end{titlepage}
\def\thefootnote{\fnsymbol{footnote}}
\setcounter{footnote}{0}

\section{Introduction and relevance of nuclear structure}

Some of my colleagues think  nuclear structure theory, in particular the nuclear shell model, is as old-fashioned as the horse-and-buggy.  
But  really it's is the exact opposite. Nuclear structure theory has lots of exciting  developments.  These developments push the shell model from 
phenomenology to rigorous first-principle calculations,  driven partly by new ideas but above all by the explosion in computing 
capabilities. While in the early days solving a $25 \times 25$ matrix was the height of computation \citep{PhysRev.105.1563}, today we find extremal eigenvalues 
of matrices exceeding dimensions of $2 \times 10^{10}$ \citep{PhysRevC.97.034328}.

Aside from the intrinsic physics interest of nuclei,  careful microscopic calculations are needed for many applications.  Detection of known and unknown particles, 
from neutrinos \citep{PhysRevC.74.034307} to dark matter \citep{PhysRevC.89.065501}, as well as experiments testing fundamental symmetries, such as neutrinoless double-$\beta$ decay 
\citep{PhysRevLett.110.222502}
and nonconservation of parity and time-reversal symmetries \citep{haxton2001atomic}, often 
require  knowledge of  matrix elements in complex nuclei.  For such calculations 
to be reliable and both precise and accurate, they need to be founded on solid microscopic calculations.  Fortunately, in many cases 
modern nuclear structure theory is rising to the challenge.

\section{Key ideas  in large-basis diagonalization}

 This paper deals solely with diagonalization of the many-body Hamiltonian in a basis built from shell-model 
single-particle states, also called the 
\textit{configuration-interaction} method or the \textit{interacting shell model} \citep{BG77,br88,ca05}.  The  idea is straightforward:  
expand a state $| \Psi \rangle$ in a basis $\{ | \alpha \rangle \}$ (assumed to be orthonormal, 
$\langle \alpha | \beta \rangle = \delta_{\alpha, \beta }$),
\begin{equation}
| \Psi \rangle = \sum_\alpha  c_\alpha  | \alpha \rangle;
\end{equation}
 minimizing $ \langle \Psi | \hat{H} | \Psi \rangle / \langle \Psi | \Psi \rangle$ leads to the eigenvalue equation
\begin{equation}
\sum_\beta H_{\alpha, \beta} c_\beta = E c_\alpha, \label{mateigen}
\end{equation}
where $H_{\alpha, \beta} = \langle \alpha | \hat{H} | \beta \rangle$ is the matrix element of the many-body Hamiltonian $\hat{H}$ in 
this basis.  I deal with the question of the choice of basis in section \ref{basis}





We can broadly classify configuration-interaction (CI) calculations into two categories, \textit{phemonenological}
 and \textit{ab initio}.  Phenomenological calcqulations are older, and usually 
assume a fixed cored and a relatively narrow valence space, such as the $1s_{1/2}$-$0d_{3/2}$-$0d_{5/2}$ space 
with a fixed $^{16}$O core, or the $1p$-$0f$ space with a fixed $^{40}$Ca core\citep{BG77,br88,ca05}.   The interactions actually start from 
some \textit{ab initio} underlying interaction, and then adjusted to many-body spectra in the target space \citep{PhysRevC.74.034315}. Because of this, 
it is fair to call them \textit{semi}-phenomenological.  By \textit{ab initio}
I mean a potential fitted to few-body data, such as nucleon-nucleon scattering and the binding energies of the $A=2,3$ and 
other light systems. These interactions are most commonly built from chiral effective field theory\citep{PhysRevC.68.041001}, 
but not always \citep{wiringa1995accurate,shirokov2016n3lo}.  
Despite having essentially the same few-body input, difference choices such as cut-off regulators \citep{PhysRevC.94.034001} can strongly influence the 
final many-body energies. 

 Purely \textit{ab initio} CI calculations are often called 
\textit{no-core shell model} (NCSM) calculations \citep{navratil2009recent,barrett2013ab}, precisely because there is no core: all particles, in principle, are active, and 
the standard methodology increases the model space until convergence: see section \ref{converge} below.

In between these two are attempts to derive \textit{ab initio} effective interactions, with no adjustable parameters, for phenomenolo\-gi\-cal-like valences spaces for medium 
and heavy nuclei, via a double projection (Okubo-Lee-Suzuki) 
method \citep{PhysRevC.91.064301}, 
via coupled clusters \citep{PhysRevLett.113.142502}, and via the 
in-medium similarity renormalization group \citep{PhysRevLett.118.032502}.

Because we cast the many-body Schr\"odinger equation as a matrix equation, the main computational task becomes solving a 
matrix eigenvalue problem.
While some bases are larger than other, as discussed below, almost all CI calculations involve large enough dimensions that it would 
be foolish to try to find all eigenpairs.  Instead, one solves for extremal eigenvalues using Arnoldi-type algorithms, almost always 
the Lanczos algorithm \citep{Lanczos}, although there have been attempts to use other methods \citep{shao2018accelerating}.

\section{Basis states for configuration interaction}

\label{basis}

How to  construct the basis set $\{ | \alpha \rangle \}$?  One choice is to use many simple states.  The most common building block 
are  Slater determinants (antisymmeterized products of single-particle states) 
or more generally the occupation-space representations of Slater determinants using creation and annihilation operators.
Furthermore, one often uses an is  \textit{M-scheme} basis, where each Slater determinant has the same fixed total 
$M$ or $J_z$, that is, the $z$-component of angular momentum. This is easy  because $J_z$ is an additive quantum number. 
Many CI shell model codes use an $M$-scheme basis, most notably {\tt ANTOINE} \citep{ANTOINE}, 
{\tt MFDn} \citep{MFDn},  {\tt BIGSTICK} 
\citep{BIGSTICK,johnson2018bigstick}, and {\tt KSHELL} \citep{shimizu2013nuclear}.
$M$-scheme bases are simple, amenable to a bit occupation representation ideal for digital computers \citep{Lanczos}, 
and one can compute matrix elements in the basis efficiently.  The drawback is one needs a large number of $M$-scheme 
basis states to build up nuclear correlations.

There are more sophisticated bases. $J$-scheme basis  states have fixed total angular momentum $J$. 
The most widely used $J$-scheme codes are {\tt OXBASH} \citep{OXBASH} and its successor {\tt NuShellX} \citep{NuShellX}.
As such, the $J$-scheme basis has smaller dimensions than the $M$-scheme. One can go even further, to so-called 
\textit{symmetry-adapted} bases, based upon groups such as SU(3) \citep{PhysRevLett.111.252501} or Sp(3,R) \citep{mccoysymplectic}. When judiciously truncated in the choice of irreps (subspaces defined by the Casimir operators of the group), such calculations can be even smaller in dimension. 

 Dimensions alone do not measure the computational burden.  From Eq.~(\ref{mateigen})  the real computational 
burden is in the nonzero matrix elements of the Hamiltonian.  $M$-scheme bases are very sparse, as small as
$\sim 10^{-6}$, while $J$-scheme bases,  smaller in dimensions, can have \textit{more} nonzero matrix elements, and 
symmetry-adapted bases yet more \citep{dytrych2016efficacy}.  Furthermore, $J$-scheme basis states are generally represented as a linear combination of 
$M$-scheme states, and symmetry-adapted states are either a linear combination of $M$-scheme states or require non-trivial recursion algorithms, making calculation of the nonzero matrix elements a significant burden; by contrast, in the $M$-scheme 
matrix elements are so simple they can be recomputed efficiently on-the-fly, dramatically reducing the memory load, albeit at a price 
of a more complicated algorithm \citep{ANTOINE,johnson2018bigstick}.  There is no `best' 
 basis,
only the recognition of trade-offs. 

In addition to the choice of many-body basis states, there is the question of the underlying single-particle basis.  Phenomenological 
calculations either assume a harmonic oscillator basis or a Woods-Saxon like basis, but in general as matrix elements are primarily 
tuned to spectra, the single-particle basis is ambiguous.  More rigorous \textit{ab initio} calculations such as the no-core shell model 
(NCSM) do have definite singe-particle bases, almost always harmonic oscillator which aids in removing spurious center-of-mass 
motion.  Yet harmonic oscillator wave functions have a steep, unphysical fall off. Hence there have been many efforts to 
introduce better wave functions \citep{PhysRevC.86.034312}, a question which has proved challenging. The most promising seem to be 
\textit{natural orbitals} \citep{constantinou2017natural}, orbitals that diagonalize the ground state one-body 
density matrix.


\section{Convergence and extrapolation}
\label{converge}

Phenomenological calculations take place in a fixed set of valence orbits (unfortunately common usage often conflates 
\textit{orbits} and \textit{shells}), with interactions tuned to that valence space, such as
the $1s$-$0d$ or $sd$-space \citep{PhysRevC.74.034315}.   \textit{Ab initio} calculations, conversely, imply a result in a unrestricted 
or infinite space.  Because any actual calculation must be done in a finite space, one must investigate the convergence as 
the space is increased, and in many cases, extrapolate to the infinite limit. 

In default NCSM calculations \citep{barrett2013ab}, one defines the model space by two parameters: the harmonic oscillator 
frequency $\Omega$, or, more often, $\hbar \Omega$, for the single-particle basis states, and $N_\mathrm{max}$, the maximum 
number of oscillator quanta allowed above the lowest configuration; historically this has also been called $N\hbar \Omega$. 
Typically one wants to extrapolate to infinite $N_\mathrm{max}$ and $\hbar \Omega$. 

One strategy is to use an exponential extrapolation, e.g. fitting energies to a form $a + b \exp (-c N_\mathrm{max})$
 \citep{heng2017ab}. This is inspired by similar exponential extrapolations in phenomenological shell model calculations 
 where even the finite model space is so large one must truncate and extrapolate  \citep{PhysRevLett.82.2064}.  
 For the NCSM, however, the results are not very robust.  Instead,
 recent work has found more robust extrapolation by combining $N_\mathrm{max}$ and $\hbar \Omega$ in to 
infrared and ultraviolet parameters, and following the convergence in 
those parameters \citep{PhysRevC.86.054002,PhysRevC.87.044326,PhysRevC.91.061301}. This can also be linked to interpreting 
$N_\mathrm{max}$ as a finite `wall' \citep{PhysRevC.86.031301}.

In a way, these extrapolations are brute force, and limited by the capability of modern computers. The basis dimension grows 
exponentially with the number of orbits / $N_\mathrm{max}$ and particles, 
which is why size-extensive methods such as coupled clusters \citep{hagen2010ab} are attractive, 
but which have their own set of limitations.   These limitations inspire  alternatives to the standard NCSM prescription: rather 
than brute force computation in a larger basis, build in \textit{smarter} bases, such as use of
better single orbitals such as natural orbitals \citep{constantinou2017natural}, and selected irreducible representations 
in symmetry-adapted bases which efficiently exploit deformation degrees of freedom \citep{PhysRevLett.111.252501,mccoysymplectic}.
These lose, however, the powerful machinery of extrapolation applied to standard NCSM calculations. 

Finally, rather than being `smarter' in our physics, one can ride a current trend and hand over insights to the computer, with
novel extrapolations  using machine learning \citep{negoita2018deep}.  The initial results are impressive, and it remains to see how 
widespread such techniques can be applied.




\Acknowledgements

This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, 
under Award Number  DE-FG02-03ER41272.

\bibliographystyle{unsrtnat}
\bibliography{johnsonmaster}




\end{document}